# ATNPA: A Unified View of Oversmoothing Alleviation in Graph Neural Networks

### v1

@@ 3.9 Transformer-based Methods
Discussion (AT):
The main difference between GraphT (Eq. 42) and GRAND (Eq. 34) is the location of the non-linearity activation function $\sigma(\cdot)$. The similarity between GRAND and transformer-based methods comes from the diffusivity modeled as an attention structure in GRAND and the diffusivity can fit into the $\texttt{Aug}(\cdot)$ Augmentation step in ATNPA. AT-type methods can be treated as a graph rewiring approach. Since the rewired graph will be more sparse compared with the original graph and the rewired process can be done in each layer, we can consider them as an extension of a controllable masking mechanism aiming to increase the energy at each layer.
@@ 4 Conclusion
In this paper, we reviewed and analyzed existing GNN oversmoothing alleviation methods. We argued that despite of dramatic differences in their design principles and math formulations, existing approaches share three common themes in their motivations to tackle oversmoothing, and such commonality allows us to summarize them into six categories. To allow in-depth understanding and analysis of all methods, we proposed ATNPA, which uses five steps to distill properties and architectures of existing methods and shows that existing oversmoothing alleviation methods are variants by introducing changes to one or multiple steps of the ATNPA. Such a unified view allows a clear understanding on how oversmoothing is alleviated for individual methods, strength and weakness of each type of methods, and possible future study directions. We drew discussion and remarks on representative methods, and observed that despite many methods focusing on constraining energy of the learned embeddings, diffusion-based methods use a physics-inspired structure to keep energy, while residual-based methods use a simple structure but focus on tuning coefficient or directly applying normalization to features to preserve energy. In addition, the modeling of network propagation has evolved from a static topology to dynamically learn topology, or to seek respective adjacency matrix for each feature instead of one shared topology for all features.

### v2

@@ 4.7 Transformer-based Methods
Discussion about AT:
The main difference between GraphT Eq. (40) and GRAND Eq. (32) is the location of the non-linearity activation function $\sigma(\cdot)$. The similarity between GRAND and transformer-based methods comes from the diffusivity modeled as an attention structure in GRAND and the diffusivity can fit into the $\texttt{Aug}(\cdot)$ Augmentation step in ATNPA. AT-type methods can be treated as a graph rewiring approach. Since the rewired graph will be more sparse compared with the original graph and the rewired process can be done in each layer, we can consider them as an extension of a controllable masking mechanism aiming to increase the energy at each layer.
5 Experimental Results Reported From Existing Methods Under Same Settings
Table 4: Collected average results from each method following the same setting for six most frequently used datasets with three homophilic and three heterophilic datasets. For the Planetoid dataset (Cora, Citeseer, Pubmed), NA indicates missing reports. Some of the results are reported by other method’s paper as its original paper has a different split setting. Bold font indicates best results and italic font indicate second best results.
Method/Dataset Year Cora Citeseer Pubmed Texas Cornell Wisconsin
APPNP 2018 83.3 71.8 80.1 65.41 73.51 69.02
GCNII 2020 85.5 73.4 80.2 77.84 76.49 81.57
GroupNorm 2020 82 69.5 79.5 NA NA NA
EGNN 2021 85.7 NA 80.1 NA NA NA
G2-gating 2023 NA NA NA 87.57 87.3 87.65
GRN 2024 NA NA NA 89.73 86.22 88.4
JKnet 2018 83.3 72.6 79.2 57.3 61.08 50.59
MixHop 2019 81.9 71.4 80.8 77.83 73.51 75.88
DAGNN 2020 84.4 73.3 80.5 NA NA NA
DropEdge 2020 81.69 71.43 79.06 NA NA NA
DropConnect 2020 82.2 71.72 NA NA NA NA
DropMessage 2023 83.33 71.83 79.2 NA NA NA
PairNorm 2020 82.1 69.6 78 60.27 58.92 48.43
NodeNorm 2020 83 72.9 80.7 78.92 80.54 83.14
WeightRep 2024 82.21 69.04 77.68 NA NA NA
GRAND 2021 84.7 73.3 80.4 NA NA NA
To verify the effectiveness of existing methods, we report each method’s best report results under similar/same settings on public datasets including three most frequently used homophilic and three heterophilic datasets. The results are shown in Table 4. For each dataset, the best performance is bold-faced, and the second best performance is in italic format. In the table, we also highlight four methods with bold fonts that achieve significant performance boost over either homophilic datasets and heterophilic datasets, including GCNII, EGNN, G2-gating, and GRN. It is worth noting that all methods are residual based with three methods focus on limiting residual coefficients based on dirichlet energy constraint or learnable metric constraint. This result suggests a promising direction following the principle of energy change regularization over propagation and residual-based GNN structure. Additionally, we observe that GRAND also shows good performance gain with its physics-informed structure, suggesting the future of dynamic system based approaches.
@@ 6 Conclusion
In this paper, we reviewed and analyzed GNN oversmoothing alleviation methods. We argued that despite of dramatic differences in their design principles and math formulations, existing approaches share three common themes in their motivations to tackle oversmoothing, and the commonality allows us to summarize them into six categories. To allow in-depth understanding and analysis of all methods, we proposed ATNPA, which uses five steps to distill properties and architectures of existing methods and shows that existing oversmoothing alleviation methods are variants by introducing changes to one or multiple steps of the ATNPA. The unified view allows a clear understanding on how oversmoothing is alleviated for individual methods, strength and weakness of each type of methods, and possible future study directions. We drew discussion and remarks on representative methods, and observed that despite many methods focusing on constraining energy of the learned embeddings, diffusion-based methods use a physics-inspired structure to keep energy, while residual-based methods use a simple structure but focus on tuning coefficient or directly applying normalization to features to preserve energy. In addition, the modeling of network propagation has evolved from a static topology to dynamically learn topology, or to seek respective adjacency matrix for each feature instead of one shared topology for all features.
The paper summarizes three main themes according to the underlying principles. With the theme dynamic system modeling combine two principles behind energy regularization and propagation and transformation decoupling. It is possible that our paper may overlook some works that are not explicitly focused on oversmoothing. Nevertheless, we are confident that reviewed methods cover major oversmoothing alleviation branches and the underling principles leveraged are coherent to the problem and are discussed to the best of our knowledge.
From a high abstract level, we highlight that oversmoothing alleviation is equivalent to dynamic system control with constraints. It is therefore suggested in the future studies to focus on two main principles for system controlling: energy initialization or boundary condition control and energy change control thorough propagation. Moreover, our ATNPA  view provides an excellent component-wise separation for future researchers to modify according to the highlighted two principles. Practically, we notice that existing methods lack a unified comparisons for large-scale benchmark on both heterophilic and homophilic datasets. Scalability of existing methods could therefore become an important future study for real-world applications.
[…]
Acknowledgements This work has been supported in part by the US National Science Foundation (NSF) under Grant Nos. IIS-2236579, IIS-2302786, and IOS-2430224.
Declarations
  • Funding: This work has been supported in part by the US National Science Foundation (NSF) under Grant Nos. IIS-2236579, IIS-2302786, and IOS-2430224.
  • Conflict of interest/Competing interests: The authors have no conflicts of interest to declare that are relevant to the content of this article.
  • Ethics approval and consent to participate: N/A
  • Consent for publication:
  • Data availability:
  • Materials availability:
  • Code availability:
  • Author contribution: Y. Jin and X. Zhu made equal contribution to the article.